# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v2

@@ D.3 Experiments on TSP & CVRP Instances with Various Distributions
Table 13: Experimental results on cross-distribution generalization. All four test sets are obtained from Omni_VRP [79] and contain 128 instances. The runtime marked with an asterisk (*) is proportionally adjusted (128/1,000) to match the size of our test datasets.
@@ D.4 Experiments on MIS Instances with Various Distributions
Table 14: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node min-max mTSP instances (100 instances). The best result is in bold.
@@
UDC-$x_{0}$($\alpha=50$) 79.5 20.10% 8s 41.0 8.62% 40s
UDC-$x_{50}$($\alpha=50$) 89.5 10.07% 9.03m 42.9 4.44% 21.05m
UDC-$x_{250}$($\alpha=50$) 94.5 5.03% 44.84m 43.8 2.41% 1.73h
@@
UDC-$x_{0}$($\alpha=20$) 21.6 23.46% 58s 116.9 12.72% 2m
UDC-$x_{50}$($\alpha=20$) 25.3 10.54% 21m 119.1 11.05% 21m
UDC-$x_{250}$($\alpha=20$) 26.4 6.59% 63m 122.9 8.25% 90m
@@
UDC-$x_{0}$($\alpha=1$) 767.0 27.09% 6s 3622.0 13.29% 2s
UDC-$x_{250}$($\alpha=1$) 901.0 14.35% 11s 3751.0 10.20% 13s
UDC-$x_{2500}$($\alpha=1$) 1009.0 4.09% 60s 4067.0 2.63% 80s
@@ D.5 Experiments on Very Large-scale
Table 15: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 5,000-node, 7,000-node CVRP (100 instances) [12, 13]. The best result is in bold and the best learning-based method is marked in shade.
@@
Table 16: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 5,000-node (128 instances), 10,000-node TSP instances (16 instances). The best result is in bold and the best learning-based method is marked in shade.
@@
Table 17: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 10,000-node, 100,000-node TSP instances (16 instances). The best result is in bold and the best learning-based method is marked in shade. We use the TSP100,000 result reported in GLOP [13] on probably a different test set.
@@ E.1 Ablation Study on $r$ (Convergence Speed In Testing)
The proposed UDC adopts the multi-conquering stage ($r$ stages) in common variants. Therefore, as an iterative method, the effectiveness of our testing phase is closely related to the testing time. We record the iterative convergence curves of the objective function value (i.e., Obj.) during testing TSP1,000, CVRP1,000, and PCTSP1,000 with UDC($\alpha$=1), UDC($\alpha$=50), GLOP [13], LEHD (RRC) [8], and DIFUSCO [39]. The results are shown in Figure 5, and curves (except for DIFUSCO, which uses reported results) are generated under the same test set and computation resources. We can obtain the following conclusions:
@@ E.2 Ablation Study on $\alpha$ (Sampled Initial Solutions)
Table 18: Ablation on the number of initial solution $r$. Objective function values (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node, and 2,000-node instances are listed.
@@ E.5 Ablation Study on Disable DCR in KP
E.6 Ablation Study on Conquering Policies
@@ Appendix F Visualization
Figure 9: Visualization of OP solutions on a random OP instance, the size of nodes from sub-figure (c) to (f) represents their values $\rho_{i}$. Obj. represents the objective function and $l(\cdot)$ is the length of solution defined in Eq. (10). The colors in subfigure (a)(b) represent sub-problems.
@@ 18. Answer: [Yes]
  19. Justification: The implementation and experiment details are presented in both the main paper (Section 4) and Appendix D and we will publish the code, datasets, and pre-trained models upon publication.
@@ 23. Answer: [Yes]
  24. Justification: We provide the code of TSP and CVRP in the supplementary file.

### v3

@@ D.3 Experiments on TSP & CVRP Instances with Various Distributions
Table 13: Experimental results on cross-distribution generalization. All four test sets are obtained from Omni_VRP [79] and contain 128 instances. The runtime marked with an asterisk (*) is proportionally adjusted (128/1,000) to match the size of our test datasets. The best learning-based method is marked by shade.
@@ D.4 Experiments on MIS Instances with Various Distributions
Table 14: Objective function (Obj.), Gap to the best algorithm (Gap) on 500-node, 1,000-node min-max mTSP instances (100 instances). For MIS, the larger objective value (Obj.) is better.
@@
UDC-$\boldsymbol{x}_{0}$($\alpha=50$) 79.5 20.10% 8s 41.0 8.62% 40s
UDC-$\boldsymbol{x}_{50}$($\alpha=50$) 89.5 10.07% 9.03m 42.9 4.44% 21.05m
UDC-$\boldsymbol{x}_{250}$($\alpha=50$) 94.5 5.03% 44.84m 43.8 2.41% 1.73h
@@
UDC-$\boldsymbol{x}_{0}$($\alpha=20$) 21.6 23.46% 58s 116.9 12.72% 2m
UDC-$\boldsymbol{x}_{50}$($\alpha=20$) 25.3 10.54% 21m 119.1 11.05% 21m
UDC-$\boldsymbol{x}_{250}$($\alpha=20$) 26.4 6.59% 63m 122.9 8.25% 90m
@@
UDC-$\boldsymbol{x}_{0}$($\alpha=1$) 767.0 27.09% 6s 3622.0 13.29% 2s
UDC-$\boldsymbol{x}_{250}$($\alpha=1$) 901.0 14.35% 11s 3751.0 10.20% 13s
UDC-$\boldsymbol{x}_{2500}$($\alpha=1$) 1009.0 4.09% 60s 4067.0 2.63% 80s
@@ D.5 Experiments on Very Large-scale
Table 15: Objective function (Obj.), Gap to the best algorithm (Gap) on 5,000-node, 7,000-node CVRP (100 instances) [12, 13]. The best result is in bold and the best learning-based method is marked in shade.
@@
Table 16: Objective function (Obj.), Gap to the best algorithm (Gap) on 5,000-node (128 instances), 10,000-node TSP instances (16 instances). The best result is in bold and the best learning-based method is marked in shade.
@@
Table 17: Objective function (Obj.), Gap to the best algorithm (Gap) on 10,000-node, 100,000-node TSP instances (16 instances). The best result is in bold and the best learning-based method is marked in shade. We use the TSP100,000 result reported in GLOP [13] on probably a different test set.
@@ E.1 Ablation Study on $r$ (Convergence Speed In Testing)
The proposed UDC adopts the multi-conquering stage ($r$ stages) in common variants. Therefore, as an iterative method, the effectiveness of our testing phase is closely related to the testing time. We record the iterative convergence curves of the objective function value (Obj.) during testing TSP1,000, CVRP1,000, and PCTSP1,000 with UDC($\alpha$=1), UDC($\alpha$=50), GLOP [13], LEHD (RRC) [8], and DIFUSCO [39]. The results are shown in Figure 5, and curves (except for DIFUSCO, which uses reported results) are generated under the same test set and computation resources. We can obtain the following conclusions:
@@ E.2 Ablation Study on $\alpha$ (Sampled Initial Solutions)
Table 18: Ablation on the number of initial solution $r$. Objective function values (Obj.), Gap to the best algorithm (Gap) on 500-node, 1,000-node, and 2,000-node instances are listed.
@@ E.5 Ablation Study on Disable DCR in KP
E.6 Supplementary Ablation Study on Conquering Policies of KP
@@ Appendix F Visualization
Figure 9: Visualization of OP solutions on a random OP instance, the size of nodes from sub-figure (c) to (f) represents their values $\rho_{i}$. Obj. represents the objective function and $l(\cdot)$ is the length of the solution defined in Eq. (10). The colors in subfigure (a)(b) represent sub-problems.
@@ 18. Answer: [Yes]
  19. Justification: The implementation and experiment details are presented in both the main paper (Section 4) and Appendix D and we have published the code, datasets, and pre-trained models via https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/UDC-Large-scale-CO-master.
@@ 23. Answer: [Yes]
  24. Justification: We provided the code of TSP and CVRP in the supplementary file for the first submission and now the final code can be found online.